In later mathematics courses, you often will want to find factors of a number with a given sum or difference. The following exercises use this technique.
Find two factors of 28 with a sum of 11. Can someone please help me with this?

Answer by brandonpark2889(31)

In later mathematics courses, you often will want to find factors of a number with a given sum or difference. The following exercises use this technique.
Find two factors of 28 with a sum of 11. Can someone please help me with this?

This can be little hard to explain. However, hear me out:
Think of all the factors of 28. If you don't know what factor means, then it is merely a a number that is multiplied with other factor to form a product.
Factors of 28:
1,2,4,7,14,28.
Now look little more carefully of the factors:
1,28
2,14
4,7
Which of the following pair has sum of 11?
1 and 28 is incorrect. 1 + 28 = 29.
2 and 14 is incorrect. 2 + 14 = 16.
4 and 7 is correct. 4 + 7 = 11.
So the answer is 4 and 7, the two factors of 28 with sum of 11.
Alright. Good luck.